OpenStudy (anonymous):

which answer shows the equation in standard form 6 - 2 x minus y equals negative 3 plus 5 A x + 2 y equals negative 1 B x + 2 y equals 1 C x - 2 y equals 1 D - 5 x + 2 y equals negative 1

OpenStudy (stamp):

6 - 2x - y = -3 + 5

OpenStudy (stamp):

6 - 2x - y = 2

OpenStudy (mimi_x3):

Ya smth looks wrong here

OpenStudy (stamp):

Subtracting six on each side of the equation gives us - 2x - y = - 4

OpenStudy (stamp):

Mimi is correct

OpenStudy (mimi_x3):

Are you sure you provided the correct question ?